1. knowing only that logo5=x and log 67=y, Find each . (3 pts) logo 35 p) logo (12)
Answer #1

(ID Given 10965 - X and log, 7=y (6) 109.6125) find (a) log 35 ca) Now 109,35 109 6(5x7) logs + log,7 I : BLOCMN) = logã +10g